In lab we have to determine how much red 40 is in the kool aid using standard addition. We plotted are data points and got y = 8.7119x+0.1037 The abosrbance of the kool aid was 1.070 and the absorbance of the red 40 was 1.058. For the kool aid we diluted .10 mL of the red 40 in 10 mL beofre putting through column to get data points. HOw do I find the cocnentration of the red 40 in the kool aid mix (0.5032 g). Then how much red 40 is in the entire kool aid packet (3g)?
Answer #1

0.1 mL in 10 mL indicates the dilution factor is 1:100  

i.e. 1.058+1.070 = 8.7119x + 0.1037

Therefore, x = The concentration of all the sample = 0.23236 M

Here, the concentration of red 40 should be greater than 0.23236 M because it is the whole concentration.

i.e. The concentration of red 40 = 0.23236*10/0.1 = 23.236 M